OpenStudy (icalibear):

What is an equation in slope-intercept form for the line that passes through the points (1,-3) and (3,1)?

OpenStudy (anonymous):

Do you know this? \[m=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (icalibear):

yea

OpenStudy (icalibear):

but for the slope i keep getting 1/2 and that doesnt work for my multiple choice options

OpenStudy (anonymous):

"m" means slope y=mx+b is the point slope form input the slope for m and any of your two coordinates for x & y, solve for b

OpenStudy (anonymous):

Then rewrite it with b & m as numbers you found, and leave the x & y as variables. That's your equation.

OpenStudy (icalibear):

the slope is 1/2 right, is that correct?

OpenStudy (icalibear):

my choices are a. y = 3x + 1 b. y = x - 3 c. 2x + 5 d. 2x - 5

OpenStudy (anonymous):

It's this... \[\frac{rise}{run}\] Take a look again at the formula I initially posted please

OpenStudy (anonymous):

3 below to 1 above is NOT 2...

OpenStudy (anonymous):

The triangle symbol, delta, in that formula means in works "the change in"

OpenStudy (icalibear):

ok well i did that and got this